# ELU-GCN: Effectively Label-Utilizing Graph Convolutional Network

### v1

@@ Appendix B Theoretical Proof
To prove Theorem 2.1, we first introduce a lemma to describe the influence of a node on the other node:
@@
The total influence is to sum over all lengths path. From Lemma B.1, we can easily obtain the influence of all labeled nodes with label $y_{1}$ on $v_{a}$ is
@@ B.2 Closed-Form Solution
$\frac{\partial\mathcal{L}}{\partial\mathbf{S}}=-2(\mathbf{Q}^{(i)})^{T}\mathbf{H}+2\mathbf{S}\mathbf{H}\mathbf{H}^{T}+2\beta\mathbf{S}$ (19)
@@
$\mathbf{S}^{(i)}=\mathbf{H}(\mathbf{Q}^{(i)})^{T}\left(\mathbf{H}\mathbf{H}^{T}+\beta\mathbf{I}_{N}\right)^{-1}.$ (20)
@@ B.3 The Woodbury identity
$=\mathbf{H}(\mathbf{Q}^{(i-1)})^{T}\left(\frac{1}{\beta}\mathbf{I}_{N}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)\mathbf{Q}^{(i-1)}.$
@@ Appendix C Related Works
This section briefly reviews the topics related to this work, including graph convolutional networks and graph structure learning.
C.1 Graph Convolutional Networks
Graph convolutional networks (GCNs) are the most popular and commonly used model in the field of graph deep learning. Early work attempted to apply the successful convolutional neural network to graph structures. For example, CheybNet (Defferrard et al., 2016) first propose that transform the graph signal from the spatial domain to the spectral domain through discrete Fourier transform, and then use polynomials to fit the filter shape (i.e., convolution). CheybNet laid the foundation for the development of spectral domain graph neural networks. The popular GCN was proposed by Kipf et al. (Kipf & Welling, 2017), which is a simplified version of ChebyNet and has demonstrated strong efficiency and effectiveness, thereby promoting the development of the graph deep learning field.
Based on the traditional GCN, many advanced GCNs have been proposed. For example, numerous works are focused on increasing the number of GCN layers. APPNP (Gasteiger et al., 2018) combines personalized random walks to expand the range of neighbors aggregated by GCN and reduce training time. JK-Net (Xu et al., 2018) integrates information from each GCN layer to enable better structure-aware representation. Additionally, there are many advanced GCNs have been proposed to increase the number of GCN layers (Chen et al., 2020; Liu et al., 2020; Wu et al., 2024). Recently, Li et al. (Li et al., 2022) developed a new framework i.e., concentration analysis, proposing a linear feature smooth method with flexible concentration properties. Huang et al. (Huang et al., 2024) found that GCN would fail on some nodes, which are often far away from the label nodes and have few neighbors, so they designed a powerful GCN model for these nodes. Liu et al. (Liu et al., 2024) proposed the graph adversarial diffusion convolution that can make GCN more robust. The core of current GCN methods is feature propagation, which allows label information to supervise the features of more nodes. However, to the best of our knowledge, no work has explored whether the label information effectively influences the features of the neighboring nodes within the GCN framework.
@@ C.2 Graph Structure Learning
Before the recent rise of Graph Neural Networks, graph structure learning had already been extensively explored from various perspectives within the field of traditional machine learning.
@@
Graph structure learning is an important technology in the graph field. It can improve the graph structure and infer new relationships between samples, thereby promoting the development of graph representation learning or other fields. Existing Graph structure learning methods can be classified into three categories, i.e., traditional unsupervised graph structure learning methods, supervised graph structure learning methods, and graph rewiring methods. Traditional unsupervised graph structure learning methods aim to directly learn a graph structure from a set of data points in an unsupervised manner. Early works (Wang & Zhang, 2006; Daitch et al., 2009) exploit the neighborhood information of each data point for graph construction by assuming that each data point can be optimally reconstructed using a linear combination of its neighbors (i.e., $\min_{A}\left\|\mathbf{AX}-\mathbf{X}\right\|_{F}^{2}$). Similarly, (Daitch et al., 2009) introduce the weight (i.e., $\min\sum_{i}\left\|\mathbf{D}_{i,i}\mathbf{X}_{i}-\sum_{j}\mathbf{A}_{i,j}\mathbf{X}_{j}\right\|^{2}$). Smoothness Jiang et al. (2019) is another widely adopted assumption on natural graph signals, the smoothness of the graph signals is usually measured by the Dirichlet energy (i.e., $\min_{\mathbf{A}}\frac{1}{2}\sum_{i,j}\mathbf{A}_{i,j}\left\|\mathbf{X}_{i}-\mathbf{X}_{j}\right\|^{2}=\min_{\mathbf{L}}\operatorname{tr}\left(\mathbf{X}^{\top}\mathbf{L}\mathbf{X}\right)$). Until now, there have been a lot of works based on the above objective function to learn graph structure. Supervised graph structure learning methods aim to use the downstream task to supervise the structure learning, which can learn a suitable structure for the downstream task. For example, NeuralSparse (Zheng et al., 2020) and PTDNet (Luo et al., 2021) directly use the adjacency matrix of the graph as a parameter and update the adjacency matrix through the downstream task. SA-SGC (Huang et al., 2023b) learns a binary classifier by distinguishing the edges connecting nodes with the same label and the edges connecting nodes with different labels in the training set, thereby deleting the edges between nodes belonging to different categories in the test set. BAGCN (Zhang et al., 2024) uses metric learning to obtain new graph structures and learns suitable metric spaces through downstream tasks. The goal of graph rewiring methods is to prevent the over-squashing (Alon & Yahav, 2021) problem. For example, FA (Alon & Yahav, 2021) proposed to use a fully connected graph as the last layer of GCN to overcome over-squashing. SDRF (Topping et al., 2022), SJLR (Giraldo et al., 2023), and BORF (Nguyen et al., 2023) aim to enhance the curvature of the neighborhood by rewiring connecting edges with small curvature. They increase local connectivity in the graph topology indirectly expanding the influence range of labels. However, the graph structure obtained by the current graph structure learning methods can not guarantee that the GNN model can effectively utilize the supervisory information transmitted in the graph.
@@ D.1 Pretaining MLP
where $\Theta^{(1)}$ and $\Theta^{(2)}$ are learnable parameters. After the above objective function converges by gradient descent algorithm, we can get $\mathbf{H}$ as follows:
@@
D.2 Training GCN with Original Graph and ELU Graph
We give details on the ELU graph-dominated GCN and the original graph-dominated GCN. In all experiments, we used a two-layer GCN. First, for the original graph-dominated GCN, we employed the vanilla GCN from:
$\overline{\mathbf{H}}=\widetilde{\mathbf{A}}\sigma(\widetilde{\mathbf{A}}\mathbf{X}\mathbf{W}^{(1)})\mathbf{W}^{(2)}$ (42)
where $\sigma(\cdot)$ is the activation function, here we use ReLU.
For the ELU graph-dominated GCN, since the $\mathbf{S}^{*}$ we obtained is actually a high-order graph adjacency matrix (i.e., $\mathbf{S}^{*}=Q^{(k)}$). Obviously, every GCN layer conducted on the $\mathbf{S}^{*}$ is non-ideally. Therefore, we use $\mathbf{S}$ in the first GCN layer and use original graph (i.e., $\widetilde{\mathbf{A}}$) in the second GCN layer:
$\widetilde{\mathbf{H}}=\widetilde{\mathbf{A}}\sigma(\mathbf{S}^{*}\mathbf{X}\mathbf{W}^{(1)})\mathbf{W}^{(2)}$ (43)
Here, to reduce model complexity, we use shared parameters in two GCNs (i.e., $\mathbf{W}^{(1)}$ and $\mathbf{W}^{(2)}$).
D.3 Details of Sparse $\mathbf{S}^{*}$ and Initialize $\mathbf{Y}$
@@
Appendix E Pseudo Code
Algorithm 2 Pseudo code of calculating $\mathbf{S}^{*}$.
0:  ELU graph $\mathbf{S}^{*}$;
1:  $\mathbf{H}=MLP(\mathbf{X})$;
2:  Expand initial labels by pseudo labels of ELU nodes;
3:  for $\text{i}\leftarrow 1,2,\cdots,k$ do
4:     Calculate $\mathbf{Q}^{(i)}$ by Eq. (LABEL:eq:10);
5:     $\mathbf{Q}^{(i)}_{l}=\mathbf{Y}_{l}$ in Eq. (LABEL:eq:10);
6:  end for
7:  Calculate $\mathbf{S}^{*}$ Eq. (9);
8:  return  $\mathbf{S}^{*}$.
Appendix F Experiments Details
F.1 Datasets
@@
Appendix G Additional Experiments
G.1 Node Classification on Social Networks
@@
G.2 Parameter Analysis
In the proposed method, we employ the non-negative parameters (i.e., $\lambda$) to achieve a trade-off between the supervised loss and the consistency loss, and $\tau$ to achieve the temperature control. To investigate the impact of $\lambda$ and $\tau$ with different settings, we conduct the node classification on the Cora and Citeseer datasets by varying the value of $\lambda$ in the range of [0.1, 1.0] and $\tau$ in the range of [0.1, 1.0]. Note that the smaller the $\tau$, the closer the model brings the positive samples and the further apart the negative samples. The results are reported in Figure 6.
From Figure 6, we have the following observations: First, the proposed method achieves significant performance when the parameter $\lambda$ or $\tau$ is in the range of [0.1, 0.2]. if $\lambda$ values are too large (e.g., >0.2) or too small (e.g., =0, the results shown in the Table 2), the performance degrades. This indicates that the proposed contrastive loss is necessary for the model. For $\tau$, setting it to 1 is equivalent to not using the temperature coefficient. The lower the temperature coefficient, the stronger the effect, indicating that $\tau$ is essential for the proposed method. Note that $\tau$ cannot be set to 0 because it is the denominator.
@@ G.2 Parameter Analysis
G.3 Analysis the Improve on $V_{\mathrm{NELU}}$
@@
Appendix H Running Time V.S. Accuracy
The biggest limitation of graph structure learning methods is the need to query in $\mathcal{O}(n^{2})$ space when learning graph structures. Although we have previously analyzed that the complexity of the proposed algorithm in graph construction is $\mathcal{O}(nc^{3})$ ($c^{3}\ll n$), plus the final graph structure is $\mathcal{O}(n^{2})$ (only one calculation is required), we further test the overall actual running time of the proposed ELU-GCN and compared with the commonly used baseline (i.e., GCN and GAT). The results are in Figure 8.
From Figure 8, we have the observations as follows. First, the overall running time of the proposed method is slightly inferior to GCN, but significantly ahead of GAT. Second, the proposed method achieves the best classification performance. Combining the above two points, the proposed method achieves the optimal trade-off between running time and model performance.
Figure 8: Running time V.S. accuracy.

### v3

@@ Appendix B Theoretical Proof
We follow the proof idea of (Wang & Leskovec, 2021), we first introduce a lemma to describe the influence of a node on the other node:
@@
The total influence is to sum over all lengths of the path. From Lemma B.1, we can easily obtain the influence of all labeled nodes with label $y_{1}$ on $v_{a}$ is
@@ B.2 Closed-Form Solution
$\frac{\partial\mathcal{L}}{\partial\mathbf{S}}=-2\mathbf{Q}^{(i)}\mathbf{H}^{T}+2\mathbf{S}\mathbf{H}\mathbf{H}^{T}+2\beta\mathbf{S}$ (21)
@@
$\mathbf{S}^{(i)}=\mathbf{Q}^{(i)}\mathbf{H}^{T}\left(\mathbf{H}\mathbf{H}^{T}+\beta\mathbf{I}_{N}\right)^{-1}.$ (22)
@@ B.3 The Woodbury identity
$=\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\frac{1}{\beta}\mathbf{I}_{N}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)\mathbf{Q}^{(i-1)}.$
@@ Appendix C Related Works
This section briefly reviews the topics related to this work, including graph convolutional networks, label propagation algorithms, and graph structure learning.
C.1 GCNs and LPA
Graph convolutional networks (GCNs) are the most popular and commonly used model in the field of graph deep learning. Early work attempted to apply the successful convolutional neural network (CNN) to graph structures. For example, CheybNet (Defferrard et al., 2016) first proposes to transform the graph signal from the spatial domain to the spectral domain through the discrete Fourier transform, and then use polynomials to fit the filter shape (i.e., convolution). CheybNet laid the foundation for the development of spectral-domain graph neural networks. The popular GCN was proposed by Kipf et al. (Kipf & Welling, 2017), which is a simplified version of ChebyNet and has demonstrated strong efficiency and effectiveness, thereby promoting the development of the graph deep learning field.
Recently, some works have focused on the combination of LPA and GCN. This is because LPA can characterize the distribution of labels spread on the graph, which can help GCN obtain more category information. Existing combined LPA and GCN methods can be classified into two categories, i.e., pseudo label methods and masked label methods. Pseudo-label methods let the output of LPA serve as the pseudo-labels to guide the representation learning. For example, PTA (Dong et al., 2021) first propagates the known labels along the graph to generate pseudo-labels for the unlabeled nodes, and second, trains normal neural network classifiers on the augmented pseudo-labeled data. GPL (Wu et al., 2024) uses the output of LPA to preserve the edges between nodes of the same class, thereby reducing the intra-class distance. The masked label methods employ LPA as regularization to assist the GCN update parameters or structures. For example, UniMP (Shi et al., 2021) makes some percentage of input label information masked at random, and then predicts it for updating parameters. GCN-LPA (Wang & Leskovec, 2021) also randomly masked a part of the labels and utilized the remaining label nodes to predict them in learning proper edge weights within labeled nodes. Although the above methods achieve excellent results on various tasks, they fail to explore under what conditions the labels propagated by LPA can best enhance the representation learning of GCNs on unlabeled nodes.
@@ C.2 Graph Structure Learning
Prior to the recent surge in Graph Neural Networks, the study of graph structure learning had already been extensively explored from multiple perspectives within traditional machine learning.
@@
Graph structure learning is an important technology in the graph field. It can improve the graph structure and infer new relationships between samples, thereby promoting the development of graph representation learning or other fields. Existing Graph structure learning methods can be classified into three categories, i.e., traditional unsupervised graph structure learning methods, supervised graph structure learning methods, and graph rewiring methods. Traditional unsupervised graph structure learning methods aim to directly learn a graph structure from a set of data points in an unsupervised manner. Early works (Wang & Zhang, 2006; Daitch et al., 2009) exploit the neighborhood information of each data point for graph construction by assuming that each data point can be optimally reconstructed using a linear combination of its neighbors (i.e., $\min_{A}\left\|\mathbf{AX}-\mathbf{X}\right\|_{F}^{2}$). Similarly, (Daitch et al., 2009) introduce the weight (i.e., $\min\sum_{i}\left\|\mathbf{D}_{i,i}\mathbf{X}_{i}-\sum_{j}\mathbf{A}_{i,j}\mathbf{X}_{j}\right\|^{2}$). Smoothness (Jiang et al., 2019) is another widely adopted assumption on natural graph signals; the smoothness of the graph signals is usually measured by the Dirichlet energy (i.e., $\min_{\mathbf{A}}\frac{1}{2}\sum_{i,j}\mathbf{A}_{i,j}\left\|\mathbf{X}_{i}-\mathbf{X}_{j}\right\|^{2}=\min_{\mathbf{L}}\operatorname{tr}\left(\mathbf{X}^{\top}\mathbf{L}\mathbf{X}\right)$). Until now, there have been a lot of works based on the above objective function to learn graph structure. Supervised graph structure learning methods aim to use the downstream task to supervise the structure learning, which can learn a suitable structure for the downstream task. For example, NeuralSparse (Zheng et al., 2020) and PTDNet (Luo et al., 2021) directly use the adjacency matrix of the graph as a parameter and update the adjacency matrix through the downstream task. SA-SGC (Huang et al., 2023b) learns a binary classifier by distinguishing the edges connecting nodes with the same label and the edges connecting nodes with different labels in the training set, thereby deleting the edges between nodes belonging to different categories in the test set. BAGCN (Zhang et al., 2024) uses metric learning to obtain new graph structures and learns suitable metric spaces through downstream tasks. The goal of graph rewiring methods is to prevent the over-squashing (Alon & Yahav, 2021) problem. For example, FA (Alon & Yahav, 2021) proposed to use a fully connected graph as the last layer of GCN to overcome over-squashing. SDRF (Topping et al., 2022), SJLR (Giraldo et al., 2023), and BORF (Nguyen et al., 2023) aim to enhance the curvature of the neighborhood by rewiring connecting edges with small curvature. They increase local connectivity in the graph topology, indirectly expanding the influence range of labels. Despite their success, existing graph structure learning methods cannot ensure that the learned graph structure effectively enables GNN models to leverage supervisory information for unlabeled nodes.
@@ D.1 Pretaining MLP
where $\Theta^{(1)}$ and $\Theta^{(2)}$ are learnable parameters. After the above objective function converges by the gradient descent algorithm, we can get $\mathbf{H}$ as follows:
@@
D.2 Details of Sparse $\mathbf{S}^{*}$ and Initialize $\mathbf{Y}$
@@
Appendix E Experiments Details
E.1 Datasets
@@
Appendix F Additional Experiments
F.1 Node Classification on Social Networks
@@
F.2 Parameter Analysis
In the proposed method, we employ the non-negative parameters (i.e., $\lambda$, $\tau$, and $eta$) to achieve a trade-off between the supervised loss and the contrastive loss, the temperature control, and the fusion of the ELU graph and the original graph. To investigate the impact of $\lambda$, $\tau$, and $\eta$ with different settings, we conduct the node classification on the Cora and Citeseer datasets by varying the value of $\lambda$, $\tau$, and $\eta$ in the range of [0.1, 1.0]. Note that the smaller the $\tau$, the closer the model brings the positive samples and the further apart the negative samples. The results are reported in Figure 5 and 6, respectively.
From Figure 5, we have the following observations: First, the proposed method achieves significant performance when the parameter $\tau$ is in the range of [0.1, 0.3] if $\tau$ values are too large (e.g., $\tau>0.3$), the performance degrades. For $\tau$, setting it to 1 is equivalent to not using the temperature coefficient. The lower the temperature coefficient, the stronger the effect, indicating that $\tau$ is essential for the proposed method. Note that $\tau$ cannot be set to 0 because it is the denominator. Second, when $\lambda$ is set in the range of [0.3, 0.6], the model can get a higher performance, if the value of $\lambda$ is too high or too small (e.g., =0, the results shown in the Table 2), the performance degrades. This indicates that the proposed contrastive loss is necessary for the model.
From Figure 6, for the parameter $\eta$, the proposed method achieves the best results while the value of the parameter is set in the range of [0.1,0.5]. This further confirms the importance of both the ELU graph and the original graph.
@@ F.2 Parameter Analysis
Figure 6: The classification performance of the proposed method at different parameter settings (i.e., $\eta$) on the Cora and Citeseer datasets.
F.3 Analysis the Improve on $V_{\mathrm{NELU}}$